Find the midspan deflection as well as the shear and moment at the support of the beam as shown. EI is constant.
W = sin x (kN/m)
L = pi (m)
1
Expert's answer
2021-10-13T14:35:41-0400

midspan deflection:


"\delta=\frac{5WL^4}{384EI}=\frac{5\pi^4sinx}{384EI}=1.27\cdot\frac{sinx}{EI}"


bending moment at any point x:

"M_x = WLx/2 -Wx2 /2=(\pi/2-x/2)xsinx"


shear:

"F=M_x/x=(\pi/2-x/2)sinx"